Considering an RCL circuit where an inductor L = 1/2 H, a resistance R = 6Ω, a capacitor C = 1/50 F and a source of e(t) = 24 sin (10t) volts are connected in series. At t = 0, the switch is closed and the current starts to flow. Considering that the capacitor has an initial voltage of 5 volts, thus,Vc (0) = 5 V and V'c (0) = 0.

While using the general form of the equation of an RCL circuit:

LC(\frac{d^2v_c{}}{dt^2}) + RC (\frac{dv_{c}}{dt}) + v_{c} = e(t)

We obtain the solution;

v_{c}(t) = e^{-6t} (\frac{75}{4}sin(8t) + 25cos(8t))-20cos(10t)

We observe that the solution has two distinct parts.

Part 1: e^{-6t} (\frac{75}{4}sin(8t) + 25cos(8t))

Part 2: -20cos(10t)

What does each part specifically represent and how are they referred to?
Answer #1

homogeneouy eqn. Soln RCL so Clotait R=64 L= 0.5H ;e= If elt) = 24 sin (lot) Ve(o) = 5V V(o)=0 - tv = e(t) RC + general en »Now 2400 sin lot Ip +12 Up + loo yp — Тоос as lot -lood sinlot - 120 C sin lot +120D coslot = 2you sinfot + louc as lot t1o